What is the y value of the y-intercept for the graph of the equation
-4x+5y=9?
Is it 9/5

Answer by AnlytcPhil(1806)


What is the y value of the y-intercept for the graph of the equation
 
-4x+5y=9? 

Is it 9/5.

Yes.

1. To find the x value of the x-intercept, substitute 0 for the other lettsr,
y, and solve for x. That will be the x value of the x intercept.

2. To find the y value of the y-intercept, substitute 0 for the other lettsr,
x, and solve for y. That will be the y value of the y intercept.

You only asked for the y value of the y intercept, so we substitute (0) for
the other letter, x,

     -4x + 5y = 9

   -4(0) + 5y = 9

      0  + 5y = 9

           5y = 9 

            y = 9/5

So you are right.  The y-intercept is the point (0, 9/5)

To find the x value of the x intercept, we substitute
(0) for the other letter, y,

     -4x + 5y = 9

   -4x + 5(0) = 9

      -4x + 0 = 9
       
          -4x = 9

            x = -9/4 
             
The x-intercept is the point (-9/4, 0)

Edwin